# A Survey of Reinforcement Learning from Human Feedback

### v1

@@ 1.4.2 PbRL and RLHF
[270] (preference-based RL).
[1] (recent advances of PbRL).
[125] (reward-rational implicit choice).
[181] (feedback types).
[45] (open issues in RLHF).
Ours (fundamentals, recent advances, and trends).
@@
Previous surveys in the domain of RLHF often focus on PbRL, where feedback is limited to binary preferences (see Section 1.2). An illustrative example of this is the survey by [270], which is a direct precursor to our work. In contrast to our work, they concentrate on binary preferences for trajectories and primarily survey methods that learn policies without deriving a reward model. Since then, the reward-modeling approach has become dominant in the field, and other approaches have extended RLHF to new feedback types. [1] give another more recent literature review of PbRL. While this review focuses on reward modeling and includes some recent work, it is far less comprehensive than our review, as many aspects are only touched upon and partly overlap with those of [270].
@@
Although not a survey per se, [125] propose reward-rational implicit choice as a unifying framework to comprehend many previous studies in PbRL and RLHF. To illustrate the generality of their approach, they overview different feedback types used in previous work and explain how they fit into their framework. Similarly, it is not strictly a survey, [181] propose a common framework for studying user interaction and interface design for multiple feedback types. As part of their work, they provide a classification of feedback types and a brief overview of RLHF approaches. Nevertheless, many facets of RLHF are not dealt with at all in those studies, as they are not primarily survey articles. In contrast to [125, 181], our survey has a more extensive coverage going beyond their study of feedback types and also discussing more recent work.
@@
[45] provide a detailed overview of the open questions and limitations of RLHF with a particular focus on aspects of security, governance, and transparency. In their article, reward modeling is also covered, as is human feedback, which goes beyond preference comparisons, but other aspects, such as theoretical approaches or an overview of existing benchmarks, are not included. Thus, it can be seen as a supplementary article that is ideal for further reading once being familiarized with the topic through our survey.
All in all, our survey can be seen as the canonical continuation of [270], which examines the evolution of the field of PbRL to the more modern and general field of RLHF. This includes a thorough description of the basics as well as an in-depth discussion of current advances and trends in the field.
@@ 1.5 Outline
In the next section, we begin with an introduction to the basics by revisiting the most important concepts from the standard RL setting, which are also naturally important in RLHF (Section 2). We then dive into the RLHF topic by outlining the most studied scenario of reward model learning from pairwise preferences. Using this introductory and illustrative example scenario, we explain the basic framework of the RLHF alongside its three main components of (human) feedback, label collection (feedback acquisition), and reward model learning. These three main components will essentially form the structure of our survey. In Section 3, we turn our attention to the human feedback component and provide an overview of the different types of feedback as well as their key attributes. The important concepts in terms of label collection are then explained in Section 4, followed by learning the reward model in Section 5. Section 6 is devoted to an overview of recent progress on the theoretical side of RLHF, including approaches involving a theoretical guarantee, as well as theoretical insights into the relationship between standard RL and RLHF. Finally, Section 7 highlights some interesting practical applications of RLHF and the existing benchmarks before Section 8 concludes the survey by pointing out some possible avenues for future work.
@@ 2 Preliminaries
In this section, we recall the basic setting and the most important concepts of RL and RLHF. In the course of this review, we will fix the notation that will be used throughout the survey. We first introduce what is probably the most studied RLHF scenario, i.e., learning a reward model from binary trajectory comparisons. On the basis of this introductory and illustrative example scenario, we explain the basic framework of RLHF with its main components and briefly discuss the respective roles of these components in the learning process. We will also briefly touch on active learning, which has a strong connection to the feedback collection component.
@@ 2.1 Reinforcement Learning
Reinforcement learning (RL) [248] is the setting of learning behavior from rewarded interaction with an environment. Such a learning environment is formalized as an Markov decision process (MDP), which is a model for sequential decision-making. In an MDP, an agent iteratively observes its current state, takes an action that causes the transition to a new state, and finally receives a reward that depends on the action’s effectiveness. Formally, an MDP is defined as a tuple $({\mathcal{S}},{\mathcal{A}},{P},{R},d_{0},\gamma)$ where
@@
Note that we here use the same notation for the return and the reward function, but both have different signatures (trajectory vs. state-action pair). We can also define return $R(\sigma)$ of a segment $\sigma$ in a similar manner. The return is well defined even if the horizon $H$ is infinite as long as $\gamma<1$. If the MDP is a tabular MDP and any trajectory has finite length, i.e., $H$ is necessarily finite, we call the MDP finite, and otherwise infinite.
ptEnvironmentptDynamicsptReward FunctionptAgentptAction $a_{t}$ptState $s_{t+1}$ptReward $r_{t+1}$
Figure 1: The standard RL setting.
@@
The basic interaction loop is depicted in Fig. 1: The agent chooses an action $a_{t}\sim\pi(s_{t})$ based on its policy and the current state. As a consequence, the environment transitions into the new state $s_{t+1}\sim P(s_{t},a_{t})$, governed by the transition dynamics. The agent observes this new state and the reward $r_{t+1}\sim R(s,a)$, after which this interaction cycle is restarted.
@@
Value-based methods (e.g., DQN and its variants [186, 115]) aim at learning the $Q$-function $Q^{*}$ of an optimal policy. The $Q$-function of a policy $\pi$ is defined by:
$Q_{\pi}(s,a)=R(s,a)+\gamma\mathbb{E}_{{P},\pi}\left[\sum_{h=0}^{H-1}\gamma^{h}{R}(s_{h},a_{h})\right]\,\text{,}$
where in the expectation, $s_{0}\sim d_{0}$, $a_{0}\sim\pi(\cdot\mid s_{0})$, and $a_{h}\sim\pi(\cdot\mid s_{h})$ as well as $s_{h}\sim{P}(\cdot\mid s_{h-1},a_{h-1})$ for $h\in[H]$. A policy can be naturally designed from a $Q$-function by choosing an action in a greedy manner in each state: $\pi(s)=\arg\max_{a}Q(s,a)$. Note that for a deterministic optimal policy $\pi^{*}$ it holds that $J(\pi^{*})=\mathbb{E}_{d_{0}}[Q^{*}(s,\pi^{*}(s))]$.
@@
$V_{\pi}(s)=\gamma\mathbb{E}_{{P},\pi}\left[\sum_{h=0}^{H-1}\gamma^{h}{R}(s_{h},a_{h})\,|\,s_{0}=s\right]\,\text{.}$
@@
In contrast, policy search methods directly aim at finding a good policy in some parametrized policy space. The most data-efficient algorithms in this class of methods follow an actor-critic scheme where both an actor (i.e., a policy) and a critic (i.e., usually its $Q$-value function) are learned at the same time. Typical representative methods here are PPO [228], TD3 [82], or SAC [103].
@@ 2.2 Preference-Based MDPs
ptEnvironmentptDynamicsptReward ModelptAgentptAction $a_{t}$ptState $s_{t+1}$ptReward $\tilde{r}_{t+1}$ptLabelerptQuery $q_{i}$ptLabel $l_{i}$
Figure 2: RLHF with reward modeling.
In contrast to standard RL as described in the previous section, RLHF does not assume that a reward signal is available. It instead assumes the existence of an oracle (e.g., human labeler) that can provide information about the reward in a specific indirect manner. More precisely, in the RLHF, the agent can make queries $q_{i}$ to the oracle, which in practice means asking for human feedback, and in response, the agent receives a label $l_{i}$, which in general gives a hint about the reward. In principle, the query can be made asynchronously to the actual conventional RL cycle. See Fig. 2 for an illustration.
In the most common setting, the oracle can compare two (segments of) trajectories, but various other cases have been considered, as we shall see later on. For the former case, RLHF is based on the setting of preference-based MDPs [94, 270], which can be defined as an MDP model without reward function, but where comparisons of trajectories are available.
@@ 2.3 Reward Learning
RLHF approaches can be divided into two categories, depending on whether a utility-based approach is used for reward modeling or an alternative criterion that is detached from a utility concept is used [95, 93, 270]. Most works fall into the first category, on which this overview focuses. Such approaches assume a human-dependent utility function that can be used as a reward function in order to apply standard RL methods. Next, we will describe the commonly used approach for reward learning for the common setting of binary trajectory comparisons.
The prevalent approach to learning a utility function from observations of pairwise comparisons is based on the Bradley-Terry model [38], which stipulates a probabilistic model for the oracle (human labeler):
@@
This entire modeling approach accommodates the case of a noisy or unreliable oracle, in which case the Bradley-Terry model can be understood as the generative model of the answers from the oracle (or labels provided by the human labeler). When the oracle is reliable, more direct methods based on preference elicitation to recover the reward function have been studied [217, 218, 264, 92, 223]. In this survey, we will focus on the general case where the oracle may be noisy.
Note that in contrast to the typical way of preference learning, the learned reward function is used to train an RL agent and not directly to compare trajectories. This discrepancy in the objective function in the reward learning part and how the learned rewards are used may lead to suboptimal policies [168].
@@ 2.4 Reinforcement Learning from Human Feedback
In the RLHF setting as illustrated in Fig. 2, the learning agent needs to solve an RL task without having access to a reward function. To this end, the agent usually simultaneously learns an approximation of the reward function (via the assumed utility function) and an RL policy. Therefore, a generic RLHF algorithm consists of repeating two phases: (1) reward learning and (2) RL training. The first phase can itself be decomposed into two main steps: (i) generate queries to ask the oracle, (ii) train a reward function approximator with the answers provided by the oracle. The RL training part is more conventional and is usually directly based on running a deep RL algorithm using the currently trained reward function approximator.
@@
This basic generic algorithm is summarized in Algorithm 1, where an off-policy actor-critic scheme is assumed to be used for the RL training part, but other RL policy learning approaches can, of course, also be used here. For an on-policy algorithm, such as PPO [228], only the recently generated transitions are used for training. For a DQN-like algorithm, lines 9 to 11 would be replaced by a loop where transitions are generated by a behavior policy based on the current estimate of the $Q$-function (e.g., $\varepsilon$-greedy algorithm) and the $Q$ network is updated using mini-batches of transitions sampled from the replay buffer $\mathcal{D}$.
@@
  • The RL agent is actually trained in a non-stationary environment since the reward approximator is concurrently updated. The RL training part needs, therefore, to account for this factor (e.g., using non-vanishing learning rates).
@@
A standard RL algorithm can be run in the RL training part, as done in most previous work in RLHF (although this may not be the best approach). This suggests that any improvements in a standard deep RL method (e.g., auxiliary losses [89], planning in learned model [109], curriculum learning [192], or data augmentation [151, 154, 166]) may potentially be transferred to the RLHF setting. In addition, most previous work in RLHF directly uses trajectories stored in replay buffer $\mathcal{D}$ to synthesize queries. An interesting research direction to explore in RLHF would be to specifically generate trajectories in order to be able to synthesize more informative queries (instead of only generating trajectories that are beneficial for RL training). This would lead to tackle a novel exploration-exploitation dilemma: Shall we visit state-action pairs that may be bad but may help better learn the reward function, or shall we visit state-action pairs that we currently think are good? This is further discussed in Section 5.5.3.
@@ 2.5 Active Learning
In active learning [230], the task is to strategically select data points for labeling to minimize the amount of labeled data required to achieve a desired level of performance, particularly valuable in scenarios like RLHF where labeling is costly. Unlike batch learning, where labeled data is predetermined, active learning empowers the learner to actively select the most informative unlabeled instances for labeling, maximizing the learning process with limited labeled data. We will only briefly introduce the active learning task here and then discuss the strategies for creating informative queries considered thus far in Section 4.
@@
At a high level, the main idea in active learning is to query data points to quickly reduce the epistemic (i.e., reducible) uncertainty about the predictions of the learned model, although other aspects can be important, such as the representativeness of the queried data points (see [261] for a survey). Two main representations are considered to describe this epistemic uncertainty: either using an ensemble of models or using a Bayesian representation. In both cases, a first basic approach selects queries using uncertainty-based criteria in order to focus on instances with high prediction uncertainty as measured by, e.g., variance or entropy computed over predictions. In contrast to the first approach, where the selection criteria are instance-based, a second approach considers criteria that may depend on all instances. Possible options are, for instance, expected model change, expected error reduction, or density-weighted uncertainty-based criteria. Here, the weights in the expectation or density allow us to take into account the distributional information about the instances and, therefore, to focus on the higher-density regions.

### v2

@@ 1.4.2 PbRL and RLHF
Wirth et al. (2017)
(preference-based RL)
Abdelkareem et al. (2022)
(recent advances of PbRL)
Jeon et al. (2020)
(feedback modelling)
Casper et al. (2023)
(open issues in RLHF)
Fernandes et al. (2023).
(language generation)
Metz et al. (2023)
(feedback types)
Yuan et al. (2024)
(feedback types)
Ours (fundamentals, recent advances, and trends)
@@
Previous surveys in the domain of RLHF often focus on PbRL, where feedback is limited to binary preferences (see Section 1.2). An illustrative example of this is the survey by Wirth et al. (2017), which is a direct precursor to our work. In contrast to our work, they concentrate on binary preferences for trajectories and primarily survey methods that learn policies without deriving a reward model. Since then, the reward-modeling approach has become dominant in the field, and other approaches have extended RLHF to new feedback types. Abdelkareem et al. (2022) give another more recent literature review of PbRL. While this review focuses on reward modeling and includes some recent work, it is far less comprehensive than our review, as many aspects are only touched upon and partly overlap with those of Wirth et al. (2017).
@@
Although not a survey per se, Jeon et al. (2020) propose reward-rational implicit choice as a unifying framework to comprehend many previous studies in PbRL and RLHF. To illustrate its generality, they overview different feedback types used in previous work and explain how they fit into their framework. The concurrent works by Metz et al. (2023) and Yuan et al. (2024), which are also not strictly surveys, propose frameworks for studying user interaction and interface design for multiple feedback types. As part of their work, they provide a classification of feedback types and a brief overview of RLHF approaches. Metz et al. (2023) have a stronger focus on the feedback interface and on learning from multiple feedback types simultaneously, discussing properties of feedback types and proposing a standard encoding for them. On the other hand, Yuan et al. (2024) also include an offline RLHF benchmark and have a stronger focus on the reward learning aspect, focusing on the entire learning pipeline. Nevertheless, many facets of RLHF are not dealt with at all in those studies, as they are not primarily survey articles. Our survey has a broader scope and, therefore, has more extensive coverage, going beyond the study of feedback types and discussing more recent work.
Domain-Specific
Fernandes et al. (2023) focuses on human feedback for langauge generation. As a result of their focus, their survey is less comprehensive than this work but discusses some language-specific aspects that do not fall into our scope, such as using feedback models at generation time.
@@
Casper et al. (2023) provide a detailed overview of the open questions and limitations of RLHF with a particular focus on aspects of security, governance, and transparency. In their article, reward modeling is also covered, as is human feedback, which goes beyond preference comparisons, but other aspects, such as theoretical approaches or an overview of existing benchmarks, are not included. Thus, it can be seen as a supplementary article that is ideal for further reading once being familiarized with the topic through our survey.
All in all, our survey can be seen as the canonical continuation of Wirth et al. (2017), which examines the evolution of the field of PbRL to the more modern and general field of RLHF. This includes a thorough description of the basics as well as an in-depth discussion of current advances and trends in the field.
@@ 1.5 Outline
In the next section, we begin with an introduction to the basics by revisiting the most important concepts from the standard RL setting, which are also naturally important in RLHF (Section 2). We then dive into the RLHF topic by outlining the most studied scenario of reward model learning from pairwise preferences. Using this introductory and illustrative example scenario, we explain the basic framework of RLHF alongside its three main components of (human) feedback, label collection (feedback acquisition), and reward model learning. These three main components will essentially form the structure of our survey. In Section 3, we turn our attention to the human feedback component and provide an overview of the different types of feedback as well as their key attributes. The important concepts in terms of label collection are then explained in Section 4, followed by learning the reward model in Section 5. We also discuss policy learning in Section 6, since some RLHF methods adapt standard RL training methods to the RLHF setting. Section 7 is devoted to an overview of recent progress on the theoretical side of RLHF, including approaches involving a theoretical guarantee, as well as theoretical insights into the relationship between standard RL and RLHF. Finally, Section 8 highlights some interesting practical applications of RLHF and the existing benchmarks before Section 9 concludes the survey by pointing out some possible avenues for future work.
@@ 2 Preliminaries
In this section, we recall the basic setting and the most important concepts of RL and RLHF. In the course of this review, we will fix the notation that will be used throughout the survey. We first introduce what is probably the most studied RLHF scenario, i.e., learning a reward model from binary trajectory comparisons. Based on this introductory and illustrative example scenario, we explain the basic framework of RLHF with its main components and briefly discuss the respective roles of these components in the learning process. We will also briefly touch on active learning, which strongly connects to the feedback collection component.
@@ 2.1 Reinforcement Learning
Reinforcement learning (RL) (Sutton & Barto, 2018) is the setting of learning behavior from rewarded interaction with an environment. Such a learning environment is formalized as an Markov decision process (MDP), which is a model for sequential decision-making. In an MDP, an agent iteratively observes its current state, takes an action that causes the transition to a new state, and finally receives a reward that depends on the action’s effectiveness. Formally, an MDP is defined as a tuple $({\mathcal{S}},{\mathcal{A}},{P},{R},d_{0},\gamma)$ where
@@
Note that we here use the same notation for the return and the reward function, but both have different signatures (trajectory vs. state-action pair). We can also define the return $R(\sigma)$ of a segment $\sigma$ in a similar manner. The return is well defined even if the horizon $H$ is infinite as long as $\gamma<1$. If the MDP is a tabular MDP and any trajectory has finite length, i.e., $H$ is necessarily finite, we call the MDP finite, and otherwise infinite.
[…]
(a) The standard RL setting.
[…]
(b) RLHF with reward modeling
Figure 1: Contrasting the standard RL setting with RLHF in its most common formulation, using a reward model. In each step, the policy commits to an action $a_{t}$ and receives the next state $s_{t+1}$ and either the true reward $r_{t+1}$ or an estimate $\hat{r}_{t+1}$ in return (symbolized by $\tilde{r}_{t+1}$). In contrast to the standard RL setting, the true reward function is not known in the RLHF setting but instead learned form human feedback. This reward learning process is decoupled from policy learning and can happen fully asynchronously. The dataset consists of a set of queries $q_{i}$ (e.g., pairs of trajectory fragments) and their labels $l_{i}$ (e.g., a preference for one of the fragments).
@@
The basic interaction loop is depicted in Fig. 1(a): The agent chooses an action $a_{t}\sim\pi(s_{t})$ based on its policy and the current state. As a consequence, the environment transitions into the new state $s_{t+1}\sim P(s_{t},a_{t})$, governed by the transition dynamics. The agent observes this new state and the reward $r_{t+1}\sim R(s,a)$, after which this interaction cycle is restarted.
@@
Value-based methods (e.g., DQN and its variants (Mnih et al., 2015; Hessel et al., 2018)) aim at learning the $Q$-function $Q^{*}$ of an optimal policy. The $Q$-function of a policy $\pi$ is defined by:
$Q_{\pi}(s,a)=\mathbb{E}_{{P},\pi}\left[\sum_{h=0}^{H-1}\gamma^{h}{R}(s_{h},a_{h})\right]\,\text{,}$
where $s_{0}=s$, and $a_{0}=a$ and in the expectation, $a_{h}\sim\pi(\cdot\mid s_{h})$ as well as $s_{h}\sim{P}(\cdot\mid s_{h-1},a_{h-1})$ for $h\in[H-1]$. A policy can be naturally designed from a $Q$-function by choosing an action in a greedy manner in each state: $\pi(s)=\arg\max_{a}Q(s,a)$. Note that for a deterministic optimal policy $\pi^{*}$ it holds that $J(\pi^{*})=\mathbb{E}_{d_{0}}[Q^{*}(s,\pi^{*}(s))]$.
@@
$V_{\pi}(s)=\mathbb{E}_{{P},\pi}\left[\sum_{h=0}^{H-1}\gamma^{h}{R}(s_{h},a_{h})\,|\,s_{0}=s\right]\,\text{.}$
@@
In contrast, policy search methods directly aim at finding a good policy in some parametrized policy space. The most data-efficient algorithms in this class of methods follow an actor-critic scheme where both an actor (i.e., a policy) and a critic (i.e., usually its $Q$-value function) are learned at the same time. Typical representative methods here are PPO (Schulman et al., 2017), TD3 (Fujimoto et al., 2018), or SAC (Haarnoja et al., 2018).
@@ 2.2 Preference-Based MDPs
In contrast to standard RL as described in the previous section, RLHF does not assume that a reward signal is available. It instead assumes the existence of an oracle (e.g., human labeler) that can provide information about the reward in a specific indirect manner. More precisely, in the RLHF, the agent can make queries $q_{i}$ to the oracle, which in practice means asking for human feedback, and in response, the agent receives a label $l_{i}$, which in general gives a hint about the reward. In principle, the query can be made asynchronously to the actual conventional RL cycle. See Fig. 1(b) for an illustration.
In the most common setting, the oracle can compare two (segments of) trajectories, but various other cases have been considered, as we shall see later on. For the former case, RLHF is based on the setting of preference-based MDPs (Gilbert et al., 2017; Wirth et al., 2017), which can be defined as an MDP model without reward function, but where comparisons of trajectories are available.
@@ 2.3 Reward Learning
RLHF approaches can be divided into two categories, depending on whether a utility-based approach is used for reward modeling or an alternative criterion that is detached from a utility concept is used (Gilbert et al., 2016; Gilbert & Weng, 2016; Wirth et al., 2017). Most works fall into the first category, on which this overview focuses. Such approaches assume a human-dependent utility function that can be used as a reward function in order to apply standard RL methods. Next, we will describe the commonly used approach for reward learning for the common setting of binary trajectory comparisons.
The prevalent approach to learning a utility function from observations of pairwise comparisons is based on the Bradley-Terry model (Bradley & Terry, 1952), which stipulates a probabilistic model for the oracle (human labeler):
@@
This entire modeling approach accommodates the case of a noisy or unreliable oracle, in which case the Bradley-Terry model can be understood as the generative model of the answers from the oracle (or labels provided by the human labeler). When the oracle is reliable, more direct methods based on preference elicitation to recover the reward function have been studied (Regan & Boutilier, 2009; 2011; Weng & Zanuttini, 2013; Gilbert et al., 2015; Sadigh et al., 2017; Wilde et al., 2018). In this survey, we will focus on the general case where the oracle may be noisy.
Note that in contrast to the typical way of preference learning, the learned reward function is used to train an RL agent and not directly to compare trajectories. This discrepancy in the objective function in the reward learning part and how the learned rewards are used may lead to suboptimal policies (Lindner et al., 2021b).
@@ 2.4 Reinforcement Learning from Human Feedback
In the RLHF setting as illustrated in Fig. 1(b), the learning agent needs to solve an RL task without having access to a reward function. To this end, the agent usually simultaneously learns an approximation of the reward function (via the assumed utility function) and an RL policy. Therefore, a generic RLHF algorithm consists of repeating two phases: (1) reward learning and (2) RL training. The first phase can itself be decomposed into two main steps: (i) generate queries to ask the oracle, (ii) train a reward function approximator with the answers provided by the oracle. The RL training part is more conventional and is usually directly based on running a deep RL algorithm using the currently trained reward function approximator.
@@
This basic generic algorithm is summarized in Algorithm 1, where an off-policy actor-critic scheme is assumed to be used for the RL training part, but other RL policy learning approaches can, of course, also be used here. For an on-policy algorithm, such as PPO (Schulman et al., 2017), only the recently generated transitions are used for training. For a DQN-like algorithm, lines 9 to 11 would be replaced by a loop where transitions are generated by a behavior policy based on the current estimate of the $Q$-function (e.g., $\varepsilon$-greedy algorithm) and the $Q$ network is updated using mini-batches of transitions sampled from the replay buffer $\mathcal{D}$.
@@
  • The RL agent is actually trained in a non-stationary environment since the reward approximator is concurrently updated. The RL training part needs, therefore, to account for this factor, e.g., using non-vanishing learning rates (see Section 6).
@@
A standard RL algorithm can be run in the RL training part, as done in most previous work in RLHF (although this may not be the best approach). This suggests that any improvements in a standard deep RL method (e.g., auxiliary losses (Gelada et al., 2019), planning in learned model (Hafner et al., 2020), curriculum learning (Narvekar et al., 2020), or data augmentation (Laskin et al., 2020; Lee et al., 2020; Lin et al., 2020b)) may potentially be transferred to the RLHF setting. In addition, most previous work in RLHF directly uses trajectories stored in replay buffer $\mathcal{D}$ to synthesize queries. An interesting research direction to explore in RLHF would be to specifically generate trajectories in order to be able to synthesize more informative queries (instead of only generating trajectories that are beneficial for RL training). This would lead to tackle a novel exploration-exploitation dilemma: Shall we visit state-action pairs that may be bad but may help better learn the reward function, or shall we visit state-action pairs that we currently think are good? This is further discussed in Section 5.5.3.
@@ 2.5 Active Learning
In active learning (Settles, 2012), the task is to strategically select data points for labeling to minimize the amount of labeled data required to achieve a desired level of performance, particularly valuable in scenarios like RLHF where labeling is costly. Unlike batch learning, where labeled data is predetermined, active learning empowers the learner to actively select the most informative unlabeled instances for labeling, maximizing the learning process with limited labeled data. We will only briefly introduce the active learning task here and then discuss the strategies for creating informative queries considered thus far in Section 4.
@@
At a high level, the main idea in active learning is to query data points to quickly reduce the epistemic (i.e., reducible) uncertainty about the predictions of the learned model, although other aspects can be important, such as the representativeness of the queried data points (see Wang et al. (2023a) for a survey). Two main representations are considered to describe this epistemic uncertainty: either using an ensemble of models or using a Bayesian representation. In both cases, a first basic approach selects queries using uncertainty-based criteria in order to focus on instances with high prediction uncertainty as measured by, e.g., variance or entropy computed over predictions. In contrast to the first approach, where the selection criteria are instance-based, a second approach considers criteria that may depend on all instances. Possible options are, for instance, expected model change, expected error reduction, or density-weighted uncertainty-based criteria. Here, the weights in the expectation or density allow us to take into account the distributional information about the instances and, therefore, to focus on the higher-density regions.